if 100 waves pass through a point of medium in 20s the ratio of frequency to time period is????
 

Grade:12th pass

1 Answers

Arun
6 years ago
100 waves pass from 20 sec
Hence in 1 sec 5 waves pass
Now
Frequency = 5 Hz
And time period = 1/5 = 0.02
Now 
The ratiof of frequency to time period = 
= 5/0.02 = 125